OpenStudy (anonymous):

find the derivative of the function y= 1/x^8

OpenStudy (anonymous):

-8x^-9

OpenStudy (anonymous):

y=1/(x^8) y=x^-8 y=-8x^-9 y=-8/(x^9)

OpenStudy (anonymous):

\[x ^{-8} = -8x ^{-9} = 8/x ^{9}\]

OpenStudy (anonymous):

umm soooo which one is the right answer haha?

OpenStudy (anonymous):

yeah its -8/(x^9) i got confused too

OpenStudy (anonymous):

alright thanks!